A vector has an x component of -21.5 units and a y component of 42.0 units. Find the magnitude and direction of the vector.
units at  ° counterclockwise from the +x-axis
Answer #1

We can find the magnitude of the vector |a| from the formula

|a| = √ax^2+ay^2

= √(-21.5)^2+(42)^2 = 47.183unit


We can find the direction of the vector from the formula:

α = tan^-1 ay/ax

= tan^-1 = 42/-21.5 = 62.891

However, we must add 180°to obtain the correct answer

α = -62.891+180 = 117.109deg